please determine the slope and the y intercept
15y+2x+6=4+2x

Answer by checkley77(12844)

y=mx+b
15y+2x+6=4+2x
15y=2x-2x+4-6
15y=-2
y=-2/15
The slope (m) here is 0.
The y intercept (b)=-2/15.